OpenStudy (anonymous):

Solve the equation. 4 times the square root of x minus 5 minus 8 = 12

OpenStudy (anonymous):

\[4\sqrt{x-5}-8=12\] Is this what you are asking?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

First step is get rid of that annoying radical sign :), but how?

OpenStudy (anonymous):

by square rooting it i think

OpenStudy (anonymous):

\[4\sqrt{x-5}-8=12\] \[4\sqrt{x-5}=20\] Divide both sides by 4 \[\sqrt{x-5}=5\] Square both sides to get rid of the root \[x-5=25\] \[x=20\] Makes sense? :)

OpenStudy (anonymous):

x=30 sorry!

terenzreignz (terenzreignz):

It does make sense :)